OpenStudy (anonymous):

Please help me solve for X in terms of pi. 2pi-x=5pi/3

OpenStudy (whpalmer4):

\[2\pi -x = \frac{5}{3}\pi\] I would start by multiplying each side by 3 to get rid of the fraction. Then add \(x\) to each side and see if you can finish from there...

OpenStudy (anonymous):

when i multiply the other side by 3, would it be 2pi-x+3?

OpenStudy (anonymous):

or would it be 6pi-x? @whpalmer4

OpenStudy (whpalmer4):

\[3(2\pi-x) = 3*2\pi - 3*x = 6\pi -3x\] Treat \(\pi\) just like you would a variable, although you mistreated the variable here :-)

OpenStudy (anonymous):

would 6pi-3x be my answer or would i have to reduce the 3 and the 6?

hero (hero):

I would have just isolated the x first.

OpenStudy (whpalmer4):

I would also, but my experience with OS users is that one can't count on their fraction manipulation skills

hero (hero):

\(2\pi -x = \frac{5}{3}\pi\) \(2\pi - \frac{5}{3} = x\) I'm confident with @ali1029 skills. I've seen it on display

hero (hero):

Ugh, forgot the \(\pi\)

OpenStudy (anonymous):

does it make a difference if the problem is actually \[2\pi-x=\frac{5\pi}{3}\] ?

hero (hero):

\(2\pi - \frac{5\pi}{3} = x\)

hero (hero):

\(2\pi = \frac{6\pi}{3}\) So \(\frac{6\pi}{3} - \frac{5\pi}{3} = x\)

OpenStudy (anonymous):

The given problem is \[2\pi-x=\frac{5\pi}{3}\]

OpenStudy (whpalmer4):

@ali1029 no, it makes no difference: \[\frac{5}{3}\pi = \frac{5\pi}{3}\]

OpenStudy (anonymous):

oh okay, just wanted to make sure! :p

OpenStudy (anonymous):

so wouldd the final answer be 6pi-3x or would i turn that into a fraction?

hero (hero):

@ali1029, did you miss what I wrote above?

OpenStudy (anonymous):

I'm sorry, let me check!

OpenStudy (anonymous):

so x = pi/3?

OpenStudy (whpalmer4):

\[2\pi -x = \frac{5\pi}{3}\]\[6\pi - 3x = 5\pi\]\[6\pi-5\pi = 3x\]\[\pi = 3x\]\[x=\frac{\pi}{3}\]

hero (hero):

Correct.

OpenStudy (anonymous):

thank you both for your help!! :)

hero (hero):

If you see an easier way to isolate x, you should do so, then simplify from there.